# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

### v2

@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
for every $k\leq\frac{g-1}{2}$. So if $g$ is odd and $k=\frac{g-1}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-1}))=0$, hence for $k>\frac{g-1}{2}$, $\mu_{2k}\equiv 0$. If $g$ is even, and $k=\frac{g-2}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-2}))=0$, hence for $k>\frac{g-2}{2}$, $\mu_{2k}\equiv 0$.
@@ 4. Schiffer variations and Weierstrass points
Let $C$ be a smooth curve of genus $g>1$, take a point $p\in C$ and fix a local coordinate $z$ centred in $p$. For $31\leq n\leq 3g-$ 3, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is given by ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $31\leq n\leq 3g-$ 3. Consider the exact sequence
@@
By Riemann Roch, if $n<2g-2$, or $n\leq 3g-3$ and $p$ is a general point, we have:
@@
Consider the hyperelliptic curve $C$ with equation (29). Fix the Weierstrass point $p=(0,0)$. For the computations of the next section it is more convenient to use the basis of $H^{0}(M)$ given by $\omega_{i}:=\alpha_{g-i}=\frac{x^{g-i}dx}{y}$, $i=1,...,g-1$. Notice that we have
@@
$=\sum\limits_{1\leq k<h\leq g-1}b_{k,h}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k}),$
@@ 5. Second fundamental form of the Hyperelliptic Torelli locus
At a point $[C]\in{\mathcal{H}}_{g}$, the dual of (42) is
@@ 6. Isotropic subspaces
To do this we will use [9, Remark 3.2], that we now recall for the reader’s convenience. Assume $Q$ is quadric in $I_{2}$, $Q=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$ where $\gamma_{1},...,\gamma_{u}$ are elements of $H^{0}(K_{C})$. Choose a local coordinate $z$ around $p$ and take a local expression of $\gamma_{\alpha}$ around $p$: $\gamma_{\alpha}=g_{\alpha}(z)dz$.
@@
([9, Remark 3.2]) Assume
@@
for $11\leq i<j\leq g-$ 1, where $L$ is the $g^{1}_{2}$, $H^{0}(L)=\langle s,t\rangle$, and $H^{0}(M)=H^{0}(K_{C}\otimes L^{\vee})=\langle\omega_{1},...,\omega_{g-1}\rangle$, where the sections $\omega_{i}$ are as in Remark 4.2. Let $Q$ be a quadric in $I_{2}$, then
$Q=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$
@@
$\mu_{2k+2}(Q)=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}=\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}\wedge\omega_{j}))=$
$=(s^{\prime}t-t^{\prime}s)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(r)}\omega_{j}^{(s)}-\omega_{i}^{(s)}\omega_{j}^{(r))})(dz)^{2k+4}\quad\forall h+l=2k+2,\quad\forall r+s=2k+1.$
@@
because $(s^{\prime}t-t^{\prime}s)(p)=0$, since $p$ is a Weierstrass point.
@@
Notice that this finishes the proof. Indeed, putting $r=g-j,m=g-i$ we have
@@
We also get $m\geq g-k$, because if $m\leq g-k-1$, $r<m$ implies $r+m<2g-2k-2$.
Recall that in Lemma 3.6 we have shown that $Q\in Ker(\mu_{2})$ if and only if for every $33\leq l\leq 2g-$ 3,
@@
By Lemma 3.7, $Q\in Ker(\mu_{2k})$, $k\geq 2$, if and only if $Q\in Ker(\mu_{2})$ and and for every $m,l$ such that $2\leq m\leq k$, $32m-1\leq l\leq 2g-$ 3,
@@
So let us prove (47). The case $k=1$ is trivial, since $42k+2=$ 4 and we only have $l=3,4$. From equations (48) we have $a_{12}=a_{13}=0$. Now assume that the result is true for $k-1$, where $k\geq 2$. Hence if $Q\in Ker(\mu_{2(k-1)})$, we have $a_{ij}=0\quad\forall l=i+j\leq 2k$. Now if $Q\in Ker(\mu_{2k})$ then $Q\in Ker(\mu_{2k-2})$, hence $a_{ij}=0\quad\forall l\leq 2k$. So we only have to check the cases $l=2k+1,2k+2$.
  • If $l=2k+1$, then we claim that $i\leq k$. Otherwise, if $i\geq k+1$, we have $j>i\geq k+1$, hence $i+j\geq 2k+3$, a contradiction. In the equations (49) with $l=2k+1$, we have $a_{1,2k},a_{2,2k-1},...,a_{k,k+1}$ as variables, since $i\leq k$. The equations given by (48) and (49) for $m=2,...,k$, $l=2k+1$, are then a system of $k$ equations in $k$ variables, which are linearly independent by the proof of Theorem 3.1. So we get $a_{i,j}=0$ for $i+j=2k+1$.
@@
Let $\omega_{1},...,\omega_{g-1}$ be our basis of $H^{0}(M)$ and let $z$ be a local coordinate centered in the Weierstrass point $p$. Assume that we locally have $\omega_{i}=g_{i}(z)dz\otimes l^{-1}$, by abuse of notation we will write $\omega_{i}^{(h)}(p)$ instead of $g_{i}^{(h)}(0)$. Then by Remark 4.2 we have
@@
We may assume $h>l$, since if $h=l$ we have trivially zero. From Remark 6.3, $\omega_{j}^{(l)}(p)=0\quad\forall j$ such that $2j+3\leq 2g-l$. Moreover, since $i<j$, this also implies $\omega_{i}^{(l)}(p)=0$, so the terms in the sum with $2j+3\leq 2g-l$ are identically zero. Then we can assume that the indexes $j$ that appear in the sum are those satisfying $12j+3\geq 2g-l+$ 1, which gives $j\geq g-\frac{l+2}{2}$. Denote by $\tilde{l}:=\frac{l+2}{2}$. Hence we have
@@
$=\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)+\sum\limits_{1\leq i<g-2}b_{i,g-2}(\omega_{i}^{(h)}\omega_{g-2}^{(l)}-\omega_{i}^{(l)}\omega_{g-2}^{(h)})(p)\\ +\dots+\sum\limits_{1\leq i<g-\tilde{l}}b_{i,g-\tilde{l}}(\omega_{i}^{(h)}\omega_{g-\tilde{l}}^{(l)}-\omega_{i}^{(l)}\omega_{g-\tilde{l}}^{(h)})(p).$
Let us focus on the first term of the above sum. From Lemma 6.2, $Q\in Ker(\mu_{2k})$ implies $b_{i,g-1}=0\quad\forall g-(2k+1)\leq i\leq g-2$. Then we have
@@
Moreover, again by Remark 6.3, if $2i+3\leq 2g-h<2g-l$, then $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$. But in the above sum we have $12i+3\leq 2(g-2k-2)+3=2g-4k-$ 1. Hence, if $h\leq 4k+1$ (so $l<4k+1$), then $2g-4k-1\leq 2g-h<2g-l$, so $2i+3\leq 2g-h<2g-l$ in all summands in (50), so the sum is identically zero.
Let us now consider the general term, where $j=g-s$. From Lemma 6.2, we have $b_{i,g-s}=0$ $\forall g-(2k+2-s)\leq i<g-s$. Hence the $i$’s that actually appear in the sum are those satisfying $i\leq g-(2k+2-s)-1=g-2k-3+s$. In the same way as we did for the first term, for $h\leq 4k+3-2s$, we have $2i+3\leq 2(g-2k-3+s)\leq 2g-h$, so $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$ and the term is zero.
It remains to look at the cases when $h\geq 4k+4-2s$. Since $h+l\leq 4k+1$ by assumption, we must have $14k+4-2s\leq h+l\leq 4k+$ 1, which gives $l\leq 2s-3$. For these values of $l$, we have $\omega_{g-s}^{(l)}(p)=0$ by Remark 6.3. Hence the term is zero again, because $i<j$, so $\omega_{i}^{(l)}(p)=0$. Thus we have shown that all terms are zero. ∎
@@
Hence the term inside the square brackets of Equation 51 becomes
@@
From Lemma 6.4 we have that
@@
Now assume $h+l=4k+2$. If $n+m\leq 4k+1$ the expression is zero exactly as before, so the only case to consider is $n+m=4k+2$ and we must have $n=h,m=l$. But in this case we have trivially zero, because $(st-ts)=0$.
We are only left to the case $h+l=4k+3$. Now again, if $m+n\leq 4k+1$ we get zero and if $m+n=4k+3$ we get zero as in the previous case (this would imply again $n=h,m=l$). So we can only have $n+m=4k+2$. This implies either $n=h-1,m=l$ or $n=h,m=l-1$. In both cases, the first factor of (52) is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$ since $p$ is a Weierstrass point. Hence we have proven that (52) is zero $\forall h+l\leq 4k+3$ and this concludes the proof.
@@
where $h+l=4k+4$. If $n+m\leq 4k+1$, we know that it is zero by Lemma 6.4. If $n+m=4k+4$, then in the equation above, we have $n=h$, $m=l$, so the first factor is (up to scalar) $(st-ts)=0$. If $n+m=4k+3$, we have either $n=h-1$ and $m=l$ or $n=h$ and $m=l-1$. In both cases, the first factor is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$, because $p$ is a Weierstrass point. So we can only have $n+m=4k+2$, where $h=n$ and $m=l-2$ or $h=n-2$ and $m=l$, since if $n=h-1$ and $m=l-1$, the first factor is $(s^{\prime}t^{\prime}-t^{\prime}s^{\prime})=0$.
@@
    $n=h$ and $m=l-2$, which only works if $l\geq 2$ and gives
@@
Let us consider the first case. By Remark 6.3, we have $\omega_{j}^{(l-2)}(p)=0$ if $22j+3\leq 2g-l+$ 2, which gives $12j\leq 2g-l-$ 1. We have two subcases here: $l$ odd or $l$ even.
  • Assume $l$ is odd. We claim that we get zero. In fact, if $j\leq g-\frac{l+1}{2}$, then $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. Hence we can assume $j\geq g-\frac{l+1}{2}+1=g-\frac{l-1}{2}$. Let $j:=g-u$, hence $u\in\{1,...,\frac{l-1}{2}\}$. From Lemma 6.2 we have $b_{i,g-u}=0$ for $i\geq g-2k-2+u$, so we may assume $i\leq g-2k-3+u$.
@@
    Now let us consider $\omega_{i}^{(h)}(p)$. We know that $\omega_{i}^{(h)}(p)=0$ if $32i\leq 2g-h-$ 3, but $i\leq g-2k-3+u$ implies $2i\leq 2g-4k-6+2u$. Now, $32g-4k-6+2u\leq 2g-h-$ 3 if and only if $2u\leq 4k+3-h$, which gives $l-1\leq 4k+3-h$, which is true if $l+h=4k+4$, that is our assumption. Hence $\omega_{i}^{(h)}(p)=0$, for all $i$ and hence every term vanishes.
  • Now assume $l$ is even. If $j\leq g-\frac{l+1}{2}=g-\frac{l}{2}-\frac{1}{2}<g-\frac{l}{2}$, we get $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. So we may assume $j\geq g-\frac{l}{2}$. Setting $j=g-u$, we get $u\leq\frac{l}{2}$. As in the odd case, we may assume $i\leq g-2k-3+u$. Now $\omega_{i}^{(l-2)}(p)=0$ if $i\leq g-\frac{l}{2}-1$ and $g-2k-3+u\leq g-\frac{l}{2}-1$ if and only if $u\leq 2k+2-\frac{l}{2}$. Since $u\leq\frac{l}{2}$, we get that the last inequality is true if and only if $l\leq 2k+2$, which is the case because $l\leq 2k$. Hence $\omega_{i}^{(l-2)}(p)=0$.
@@
So let us assume $l$ even, then also $h$ is even. By Remark 6.3, $\omega_{j}^{(l-2)(p)}=0$ if $j\leq g-\frac{l}{2}-\frac{3}{2}<g-\frac{l}{2}-1$, so we may assume $j\geq g-\frac{l}{2}-1$. As usual we put $j=g-u$ and we have $u\leq\frac{l}{2}+1$. We have $\omega_{i}^{(l)}(p)=0$ if $32i\leq 2g-l-$ 3, but, as in the previous case, we can assume $i\leq g-2k-3+u$. Since $l\leq 2k$, and $u\leq\frac{l}{2}+1$, we have $32g-6-4k+2u\leq 2g-l-$ 3. So $32i\leq 2g-6-4k+2u\leq 2g-l-$ 3, and hence $\omega_{i}^{(l)}(p)=0$.
Again, by Remark 6.3, $\omega_{i}^{(h-2)}(p)=0$ if $12i\leq 2g-h-$ 1. As usual we may assume $2i\leq 2g-6-4k+2u$. We have $u\leq\frac{l}{2}+1$. We claim that the only nonzero term occurs for $u=\frac{l}{2}+1$ and $i=g-\frac{h}{2}$. In fact, if $u\leq\frac{l}{2}$, we have $12g-6-4k+2u\leq 2g-h-$ 1, so $\omega_{i}^{(h-2)}(p)=0$. So $u=\frac{l}{2}+1$, $i=g-\frac{h}{2}$ and the only term is $b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$
@@
  • $-(st^{\prime\prime}-ts^{\prime\prime})(p)\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$ if $l=0$ and $h=4k+4$.
@@
Notice that every term in the equation is a linear combination of $b_{r,u}$ where $r+u=2g-2k-3$.
Let $\lambda_{k,u}$ be the coefficient of $b_{g-2k-3+u,g-u}$ where $u=1,...,k+1$. We will now see that $\lambda_{k,u}\neq 0$ for every $k,u$. For $1\leq u\leq k$, we have
@@
$-8u^{3}+8u^{2}(k+1)-4ku-2k-3\neq 0$
because it is odd for every $k,u$. Moreover, $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)\neq 0$ because $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)$ has order $4k+4-2u$ at $p$ and $\omega_{g-u}^{(2u-2)}$ has order $22u-$ 2 at $p$ (see Lemma 4.2).

### v3

@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
for every $k\leq\frac{g-1}{2}$. So if $g$ is odd and $k=\frac{g-1}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-1}))=0$, hence for $k>\frac{g-1}{2}$, $\mu_{2k}\equiv 0$. If $g$ is even, and $k=\frac{g-2}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-2}))=0$, hence for $k>\frac{g-2}{2}$, $\mu_{2k}\equiv 0$, since the domain is 0.
@@ 4. Schiffer variations and Weierstrass points
Let $C$ be a smooth curve of genus $g>1$, take a point $p\in C$ and fix a local coordinate $z$ centred in $p$. For $1\leq n\leq 3g-3$, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is given by $\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $1\leq n\leq 3g-3$. Consider the exact sequence
@@
By Riemann Roch, if n<2g2n<2g-2, or $n\leq 3g-3$ and $p$ is a general point, we have:
@@
Consider the hyperelliptic curve $C$ with equation (20). Fix the Weierstrass point $p=(0,0)$. For the computations of the next section it is more convenient to use the basis of $H^{0}(M)$ given by $\omega_{i}:=\alpha_{g-i}=\frac{x^{g-i}dx}{y}$, i=1,,g1i=1,...,g-1. Notice that we have
@@
$\sum\limits_{1\leq k<h\leq g-1}b_{k,h}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k}),$
@@ 5. Second fundamental form of the Hyperelliptic Torelli locus
At a point $C]\in{\mathcal{H}}_{g}$, the dual of (36) is
@@ 6. Isotropic subspaces
To do this we will use [9, Remark 3.2], that we now recall for the reader’s convenience.
@@
([9, Remark 3.2]) Let $Q=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}\gamma_{\alpha}\odot\gamma_{\beta}$ be a quadric in $I_{2}$, where $\gamma_{1},...,\gamma_{u}$ are elements of $H^{0}(K_{C})$. Choose a local coordinate $z$ around $p$ and take a local expression of $\gamma_{\alpha}$ around $p$: $\gamma_{\alpha}=g_{\alpha}(z)dz$.
Assume moreover that
@@
for $1\leq i<j\leq g-1$, where $L$ is the $g^{1}_{2}$, $H^{0}(L)=\langle s,t\rangle$, and $H^{0}(M)=H^{0}(K_{C}\otimes L^{\vee})=\langle\omega_{1},...,\omega_{g-1}\rangle$, where the sections $\omega_{i}$ are as in Remark 4.2. Let $Q$ be a quadric in $I_{2}$, then
$Q=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\odot\gamma_{\beta},$
@@
$\mu_{2k+2}(Q)=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(dz)^{2k+4}=\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}\wedge\omega_{j}))(dz)^{2k+4}=$
$(f^{\prime}g-fg^{\prime})\sum\limits_{1\leq i<j\leq g-1}b_{ij}(f_{i}^{(r)}f_{j}^{(s)}-f_{i}^{(s)}f_{j}^{(r))})(dz)^{2k+4}\quad\forall h+l=2k+2,\quad\forall r+s=2k+1,$
where in a local coordinate we have $s=f(z)T$, $t=g(z)T$, $\omega_{i}=f_{i}(z)dz\otimes T^{\vee}$, $\forall i$.
@@
because $(f^{\prime}g-fg^{\prime})(p)=0$, since $p$ is a Weierstrass point.
@@
Notice that this finishes the proof. Indeed, putting r=gj,m=gir=g-j,m=g-i we have
@@
We also get $m\geq g-k$, because if $m\leq g-k-1$, $r<m$ implies r+m<2g2k2r+m<2g-2k-2.
Recall that in Lemma 3.6 we have shown that $Q\in Ker(\mu_{2})$ if and only if for every $3\leq l\leq 2g-3$,
@@
By Lemma 3.7, $Q\in Ker(\mu_{2k})$, $k\geq 2$, if and only if $Q\in Ker(\mu_{2})$ and and for every $m,l$ such that $2\leq m\leq k$, $2m-1\leq l\leq 2g-3$,
@@
So let us prove (41). The case $k=1$ is trivial, since $2k+2=4$ and we only have $l=3,4$. From equations (42) we have $a_{12}=a_{13}=0$. Now assume that the result is true for k1k-1, where $k\geq 2$. Hence if $Q\in Ker(\mu_{2(k-1)})$, we have $a_{ij}=0\quad\forall l=i+j\leq 2k$. Now if $Q\in Ker(\mu_{2k})$ then $Q\in Ker(\mu_{2k-2})$, hence $a_{ij}=0\quad\forall l\leq 2k$. So we only have to check the cases $l=2k+1,2k+2$.
  • If $l=2k+1$, then we claim that $i\leq k$. Otherwise, if $i\geq k+1$, we have $j>i\geq k+1$, hence $i+j\geq 2k+3$, a contradiction. In the equations (43) with $l=2k+1$, we have $a_{1,2k},a_{2,2k-1},...,a_{k,k+1}$ as variables, since $i\leq k$. The equations given by (42) and (43) for m=2,,km=2,...,k, $l=2k+1$, are then a system of $k$ equations in $k$ variables, which are linearly independent by the proof of Theorem 3.1. So we get $a_{i,j}=0$ for $i+j=2k+1$.
@@
Let $\omega_{1},...,\omega_{g-1}$ be our basis of $H^{0}(M)$ and let $z$ be a local coordinate centered in the Weierstrass point $p$. Assume that we locally have $\omega_{i}=f_{i}(z)dz\otimes T^{\vee}$, by abuse of notation we will write $\omega_{i}^{(h)}(p)$ instead of $f_{i}^{(h)}(0)$. Then by Remark 4.2 we have
@@
We may assume $h>l$, since if $h=l$ we have trivially zero. From Remark 6.3, $\omega_{j}^{(l)}(p)=0\quad\forall j$ such that $2j+3\leq 2g-l$. Moreover, since $i<j$, this also implies $\omega_{i}^{(l)}(p)=0$, so the terms in the sum with $2j+3\leq 2g-l$ are identically zero. Then we can assume that the indexes $j$ that appear in the sum are those satisfying $2j+3\geq 2g-l+1$, which gives $j\geq g-\frac{l+2}{2}$. Denote by $\tilde{l}:=\frac{l+2}{2}$. Hence we have
@@
$\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)+\sum\limits_{1\leq i<g-2}b_{i,g-2}(\omega_{i}^{(h)}\omega_{g-2}^{(l)}-\omega_{i}^{(l)}\omega_{g-2}^{(h)})(p)\\ +\dots+\sum\limits_{1\leq i<g-\tilde{l}}b_{i,g-\tilde{l}}(\omega_{i}^{(h)}\omega_{g-\tilde{l}}^{(l)}-\omega_{i}^{(l)}\omega_{g-\tilde{l}}^{(h)})(p).$
Let us focus on the first term of the above sum. From Lemma˜6.2, $Q\in Ker(\mu_{2k})$ implies $b_{i,g-1}=0\quad\forall g-(2k+1)\leq i\leq g-2$. Then we have
@@
Moreover, again by Remark˜6.3, if $2i+3\leq 2g-h<2g-l$, then $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$. But in the above sum we have $2i+3\leq 2(g-2k-2)+3=2g-4k-1$. Hence, if $h\leq 4k+1$ (so $l<4k+1$), then $2g-4k-1\leq 2g-h<2g-l$, so $2i+3\leq 2g-h<2g-l$ in all summands in (44), so the sum is identically zero.
Let us now consider the general term, where j=gsj=g-s. From Lemma˜6.2, we have $b_{i,g-s}=0$ $\forall g-(2k+2-s)\leq i<g-s$. Hence the $i$’s that actually appear in the sum are those satisfying $i\leq g-(2k+2-s)-1=g-2k-3+s$. In the same way as we did for the first term, for $h\leq 4k+3-2s$, we have $2i+3\leq 2(g-2k-3+s)\leq 2g-h$, so $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$ and the term is zero.
It remains to look at the cases when $h\geq 4k+4-2s$. Since $h+l\leq 4k+1$ by assumption, we must have $4k+4-2s\leq h+l\leq 4k+1$, which gives $l\leq 2s-3$. For these values of $l$, we have $\omega_{g-s}^{(l)}(p)=0$ by Remark 6.3. Hence the term is zero again, because $i<j$, so $\omega_{i}^{(l)}(p)=0$. Thus we have shown that all terms are zero. ∎
@@
Hence the term inside the square brackets of Equation˜45 becomes
@@
From Lemma˜6.4 we have that
@@
Now assume $h+l=4k+2$. If $n+m\leq 4k+1$ the expression is zero exactly as before, so the only case to consider is $n+m=4k+2$ and we must have $n=h,m=l$. But in this case we have trivially zero, because (stts)=0(st-ts)=0.
We are only left to the case $h+l=4k+3$. Now again, if $m+n\leq 4k+1$ we get zero and if $m+n=4k+3$ we get zero as in the previous case (this would imply again $n=h,m=l$). So we can only have $n+m=4k+2$. This implies either n=h1,m=ln=h-1,m=l or n=h,m=l1n=h,m=l-1. In both cases, the first factor of (46) is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$ since $p$ is a Weierstrass point. Hence we have proven that (46) is zero $\forall h+l\leq 4k+3$ and this concludes the proof.
@@
where $h+l=4k+4$. If $n+m\leq 4k+1$, we know that it is zero by Lemma˜6.4. If $n+m=4k+4$, then in the equation above, we have $n=h$, $m=l$, so the first factor is (up to scalar) (stts)=0(st-ts)=0. If $n+m=4k+3$, we have either n=h1n=h-1 and $m=l$ or $n=h$ and m=l1m=l-1. In both cases, the first factor is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$, because $p$ is a Weierstrass point. So we can only have $n+m=4k+2$, where $h=n$ and m=l2m=l-2 or h=n2h=n-2 and $m=l$, since if n=h1n=h-1 and m=l1m=l-1, the first factor is $(s^{\prime}t^{\prime}-t^{\prime}s^{\prime})=0$.
@@
    $n=h$ and m=l2m=l-2, which only works if $l\geq 2$ and gives
@@
Let us consider the first case. By Remark 6.3, we have $\omega_{j}^{(l-2)}(p)=0$ if $2j+3\leq 2g-l+2$, which gives $2j\leq 2g-l-1$. We have two subcases here: $l$ odd or $l$ even.
  • Assume $l$ is odd. We claim that we get zero. In fact, if $j\leq g-\frac{l+1}{2}$, then $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. Hence we can assume $j\geq g-\frac{l+1}{2}+1=g-\frac{l-1}{2}$. Let j:=guj:=g-u, hence $u\in\{1,...,\frac{l-1}{2}\}$. From Lemma˜6.2 we have $b_{i,g-u}=0$ for $i\geq g-2k-2+u$, so we may assume $i\leq g-2k-3+u$.
@@
    Now let us consider $\omega_{i}^{(h)}(p)$. We know that $\omega_{i}^{(h)}(p)=0$ if $2i\leq 2g-h-3$, but $i\leq g-2k-3+u$ implies $2i\leq 2g-4k-6+2u$. Now, $2g-4k-6+2u\leq 2g-h-3$ if and only if $2u\leq 4k+3-h$, which gives $l-1\leq 4k+3-h$, which is true if $l+h=4k+4$, that is our assumption. Hence $\omega_{i}^{(h)}(p)=0$, for all $i$ and hence every term vanishes.
  • Now assume $l$ is even. If $j\leq g-\frac{l+1}{2}=g-\frac{l}{2}-\frac{1}{2}<g-\frac{l}{2}$, we get $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. So we may assume $j\geq g-\frac{l}{2}$. Setting j=guj=g-u, we get $u\leq\frac{l}{2}$. As in the odd case, we may assume $i\leq g-2k-3+u$. Now $\omega_{i}^{(l-2)}(p)=0$ if $i\leq g-\frac{l}{2}-1$ and $g-2k-3+u\leq g-\frac{l}{2}-1$ if and only if $u\leq 2k+2-\frac{l}{2}$. Since $u\leq\frac{l}{2}$, we get that the last inequality is true if and only if $l\leq 2k+2$, which is the case because $l\leq 2k$. Hence $\omega_{i}^{(l-2)}(p)=0$.
@@
So let us assume $l$ even, then also $h$ is even. By Remark 6.3, $\omega_{j}^{(l-2)(p)}=0$ if $j\leq g-\frac{l}{2}-\frac{3}{2}<g-\frac{l}{2}-1$, so we may assume $j\geq g-\frac{l}{2}-1$. As usual we put j=guj=g-u and we have $u\leq\frac{l}{2}+1$. We have $\omega_{i}^{(l)}(p)=0$ if $2i\leq 2g-l-3$, but, as in the previous case, we can assume $i\leq g-2k-3+u$. Since $l\leq 2k$, and $u\leq\frac{l}{2}+1$, we have $2g-6-4k+2u\leq 2g-l-3$. So $2i\leq 2g-6-4k+2u\leq 2g-l-3$, and hence $\omega_{i}^{(l)}(p)=0$.
Again, by Remark 6.3, $\omega_{i}^{(h-2)}(p)=0$ if $2i\leq 2g-h-1$. As usual we may assume $2i\leq 2g-6-4k+2u$. We have $u\leq\frac{l}{2}+1$. We claim that the only nonzero term occurs for $u=\frac{l}{2}+1$ and $i=g-\frac{h}{2}$. In fact, if $u\leq\frac{l}{2}$, we have $2g-6-4k+2u\leq 2g-h-1$, so $\omega_{i}^{(h-2)}(p)=0$. So $u=\frac{l}{2}+1$, $i=g-\frac{h}{2}$ and the only term is $b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$
@@
  • $(st^{\prime\prime}-ts^{\prime\prime})(p)\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$ if $l=0$ and $h=4k+4$.
@@
Notice that every term in the equation is a linear combination of $b_{r,u}$ where r+u=2g2k3r+u=2g-2k-3.
Let $\lambda_{k,u}$ be the coefficient of $b_{g-2k-3+u,g-u}$ where u=1,,k+1u=1,...,k+1. We will now see that $\lambda_{k,u}\neq 0$ for every $k,u$. For $1\leq u\leq k$, we have
@@
$8u^{3}+8u^{2}(k+1)-4ku-2k-3\neq 0$
because it is odd for every $k,u$. Moreover, $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)\neq 0$ because $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)$ has order 4k+42u4k+4-2u at $p$ and $\omega_{g-u}^{(2u-2)}$ has order 2u22u-2 at $p$ (see Lemma 4.2).